# Perturbations of relativistic dissipative stars

### v1

@@ Abstract
Viscous fluids can dissipate and alter the propagation of gravitational waves, as well as modify the relaxation and stability properties of self-gravitating fluids. This is particularly relevant in order to understand the relaxation to equilibrium of neutron stars, and their gravitational wave emission. Here we study the linearized theory of perturbations of spherically symmetric self-gravitating fluids. Dissipative effects are included through the hydrodynamics theory of Bemfica, Disconzi, Noronha and Kovtun (BDNK). This theory has been shown to be causal and stable, despite involving only first order gradients. We show how the problem reduces to two coupled wave equations in the axial sector, one of them associated to a novel viscous mode, and including explicitly dissipative terms. In the polar sector we reduce the problem to five coupled wave equations and one additional constraint. We comment on their causal structure, and recover the causality bounds of the BDNK theory.
  • Niels Bohr International Academy, Niels Bohr Institute, Blegdamsvej 17, 2100 Copenhagen, Denmark
@@ 1 Introduction
Understanding the interaction of GWs and viscous matter also becomes important in the context of precision GW astrophysics. GW interferometers have and will observe mergers of neutron stars (NSs), which then might relax to equilibrium, or collapse promptly to form a black hole. Viscous effects can affect their GW signal during the inspiral stage Hegade K. R. et al. (2023); Ripley et al. (2023a, b); Chia et al. (2024), but also the GW emission during their final relaxation. Numerical simulations hint towards a significant impact, in particular, of bulk viscosity, in the late time relaxation of NSs Alford et al. (2010, 2018); Alford and Harris (2018); Most et al. (2021, 2022); Camelio et al. (2023a, b); Chabanov and Rezzolla (2023a, b). This happens because newly formed NSs are very hot, and bulk viscosity becomes the dominant dissipative effect at high temperatures. However, the only perturbative estimates on the relaxation timescales due to dissipative effects Cutler and Lindblom (1987); Cutler et al. (1990) are, indeed, estimates, based on the energy functional of Lindblom and Detweiler (1983). A first-principles calculation of the impact of different kinds of viscosity on the quasinormal modes of hot NSs, their precise damping times, and so on, is still lacking.
@@
In order to do this, we need to specify a particular theory of dissipative hydrodynamics. The naive extensions of the Navier–Stokes hydrodynamics to the relativistic regime, first undertaken by Eckart Eckart (1940) and Landau Landau and Lifshitz (2013) are problematic: they lead to instabilities because of their acausal character Hiscock and Lindblom (1983, 1985); Detweiler and Lindblom (1985). There is a way to heal these pathologies, by extending the theory and including effects which are second order in the gradients. This is sometimes referred to as the Israel–Stewart theory Israel and Stewart (1979); Rezzolla and Zanotti (2013), or as extended irreversible thermodynamics. However, the inclusion of second order effects complicates significantly the analysis. Recent work tackled the problem of linearized perturbations in this theory Díaz-Guerra et al. (2024). An alternative approach, based on a variational framework, was put forward by Carter Carter (1985, 1988, 2006); Andersson and Comer (2021). Perhaps the most natural way to solve the problems that the Eckart or Landau theories suffer from is solved simply by including all possible first order corrections to the stress energy tensor. This means including also timelike gradients, and not just spacelike ones. By following this line of thought, Bemfica, Disconzi, Noronha and Kovtun (BDNK) Bemfica et al. (2018, 2019a, 2019b, 2022); Kovtun (2019) showed that one can obtain a theory that is (i) causal, provided the transport coefficients satisfy some constraints, (ii) linearly stable, and (iii) the stress energy tensor only involves up to first order covariant derivatives of the primary thermodynamical variables. For this reason, in this work we will consider the stress energy tensor of a fluid in the BDNK theory as a starting point.
@@
$c_{s}$ $\sqrt{\partial\mathrm{p}/\partial\mathrm{e}}$ (Sound speed)  (17)
@@ 2 GSGM Formalism for spacetime perturbations
We use spherical symmetry to split the spacetime manifold as a warped product of a sphere and a two dimensional Lorentzian manifold, i.e., $\mathcal{M}=\mathcal{M}^{2}\times_{r^{2}}\mathcal{S}^{2}$. This decomposition allows us to reduce Einstein equations, and its perturbations, to equations on $\mathcal{M}^{2}$. Gerlach and Sengupta first used this idea to obtain a particularly elegant derivation of the equations describing linearized perturbations around a Schwarzschild black hole Gerlach and Sengupta (1979a, 1980). Since then, this formalism has been later revisited and adapted to different set-ups, ranging from higher dimensions Kodama et al. (2000); Ishibashi and Kodama (2003); Kodama and Ishibashi (2003, 2004), further restrictions to vacuum spacetimes Martel and Poisson (2005), including deriving the equations from the curvature wave equation Mukkamala and Pereñiguez (2024), higher order perturbations Brizuela et al. (2006, 2007), perturbations in modified gravity theories Ripley and Yagi (2018); Tattersall et al. (2018), gauge theories Pereñiguez (2023), and non-stationary spacetimes Redondo-Yuste et al. (2024).
@@ 2.1 Background geometry
where $r$ is the warping factor (areal radius) and $\gamma_{ab}$ is the usual metric on the $2$–sphere. Let us introduce $v_{A}=\nabla_{A}r$. Here, $\nabla$ denotes covariant derivatives on $\mathcal{M}^{2}$ compatible with the metric $g_{AB}$, and $D$ the covariant derivative on the sphere.
@@ 2.2 Linear perturbations
We now consider linearized perturbations of the metric and the stress–energy tensor. We expand them in tensor spherical harmonics following Martin-Garcia and Gundlach (1999, 2001). The tensor spherical harmonics $\{\mathcal{Z}^{\ell},\mathcal{X}^{\ell}\}$ are introduced in the Supplemental Material (SM). We also write already the metric perturbation in terms of gauge invariant variables (equivalently, in the Regge-Wheeler gauge), and discuss gauge invariance in the SM. The metric and stress energy tensor perturbations can be written as
@@
$+2\Bigl{(}k^{\ell}_{n}n_{A}-k^{\ell}_{l}l_{A}\Bigr{)}\mathcal{X}^{\ell}_{a}dy^{A}d\theta^{a}+r^{2}h^{\ell}_{S}\gamma_{ab}\mathcal{Z}^{\ell}\gamma_{ab}d\theta^{a}d\theta^{b}\,,$
@@
$+\Bigl{(}4\pi(t_{g}-t_{p}+4t_{S})+2(W^{2}+\mu U-\nu W)-v^{2}-\nu^{\prime}+\frac{\lambda^{2}-1}{r^{2}}\Bigr{)}k_{l}=16\pi\vartheta_{l}\,,$
@@
$+\Bigl{(}4\pi(2t_{g}+t_{p}+2t_{S})+2(\mu U-\nu W-W^{2})+\nu^{\prime}+\nu^{2}-\mu^{2}+\frac{\lambda^{2}}{r}\Bigr{)}k_{n}=16\pi\vartheta_{n}\,,$
@@
$\mathcal{E}_{l}\equiv$ $h^{\prime}_{q}-\dot{h}_{g}-\dot{h}_{p}-2\dot{h}_{S}+2\Bigl{(}Uh_{g}+\nu h_{q}-\mu h_{P})=32\pi\mathcal{T}_{l}\,,$
$\mathcal{E}_{n}\equiv$ $h^{\prime}_{p}-h^{\prime}_{g}-\dot{h}_{q}-2h^{\prime}_{S}+2\Bigl{(}Wh_{g}+\nu h_{p}-\mu h_{q}\Bigr{)}=32\pi\mathcal{T}_{n}\,,$
~$\tilde{\mathcal{E}}_{S}\equiv$ $h_{g}=-16\pi r^{2}\tilde{\mathcal{T}}_{S}\,,$
@@
$+2\Bigl{[}2\nu(\mu+U)+2\mu W+\dot{\nu}-\mu^{\prime}-4\pi t_{q}\Bigr{]}h_{q}-\Bigl{(}16\pi t_{S}+\frac{\lambda^{2}}{r^{2}}\Bigr{)}h_{g}+32\pi t_{S}h_{S}=32\pi\mathcal{T}_{S}\,,$
@@ 3.1 The general theory
$T_{\mu\nu}=\mathcal{E}u_{\mu}u_{\nu}+\mathcal{P}\perp_{\mu\nu}+2u_{(\mu}\mathcal{Q}_{\nu)}+\mathcal{T}_{\mu\nu}\,,$ (16)
@@
The quantities $\mathcal{E},\mathcal{P},\mathcal{Q},\mathcal{T}$ need to be specified as functions of some primary variables. A general fluid is described by three primary variables: the fluid velocity $u^{\mu}$, its energy density $\mathrm{e}$ and its particle number density. In this work, in order to simplify the problem, we will consider that heat conductivity is negligible, and hence we restrict to barotropic matter. We leave for future explorations the inclusion of heat diffusion, and a chemical potential.
Now, specifying how do the terms of the stress energy tensor depend on the primary variables $\{u^{\mu},\mathrm{e}\}$ characterizes the fluid theory. For a perfect fluid, for example, one just requires that $\mathcal{E}=\mathrm{e}$, $\mathcal{P}=\mathrm{p}$ is the pressure, which is related to the energy by some equation of state, and $\mathcal{Q}=\mathcal{T}=0$. For BDNK fluids we consider additionally terms involving first order spacetime derivatives of the primary variables. In detail, the constitutive relations are
@@
where we have introduced the shorthand $\rho=\mathrm{e}+\mathrm{p}$, $\sigma_{\mu\nu}$ is the shear associated to $u_{\mu}$, and $\{\tau_{\mathrm{e}},\tau_{\mathrm{p}},\tau_{\mathcal{Q}},\eta,\zeta\}$ are the transport coefficients. Notice that the transport coefficients $\tau_{\mathrm{e},\mathrm{p},\mathcal{Q}}$ are effectively relaxation timescales, which are necessary to be added to make the theory causal. We have also introduced the sound speed $c_{s}^{2}=\partial\mathrm{p}/\partial\mathrm{e}$. We will also introduce the following combinations of transport coefficients
@@
The transport coefficients are obtained from the microscopic theory. In order for the resulting hydrodynamic theory to be causal and stable, they must satisfy certain inequalities Bemfica et al. (2022). For example, in a companion work Boyanov et al. (2024) we choose a particular dependence of the transport coefficients on the stellar structure, which in turn constraints the maximum shear viscosity. However, in order to be as general as possible, here we just assume that they are smooth functions of the radial coordinate alone (in particular, we assume that they do not vary with time).
From these constitutive relations, we can immediately obtain the components of the stress energy tensor of the background. Spherical symmetry requires the fluid velocity to be $u_{\mu}dx^{\mu}=u_{A}dy^{A}$. As is natural, we choose a frame where $l_{A}=u_{A}$, and the normal vector $n_{A}=\epsilon_{AB}u^{B}$. Then, the constitutive relations become
@@ 3.2 Fluid perturbations
Now, in terms of these perturbed variables, we construct the stress energy tensor, and evaluate each of its components. A straightforward but lengthy calculation yields
@@ 4 Coordinates
The previous expressions are valid for any coordinate chart (for the even parity case, for any stationary coordinate chart), which has the same domain of definition as the frame that we have introduced. We highlight that a frame is only locally defined and might not be defined, e.g., both in the exterior and in the interior region of a horizon, when dealing with black hole spacetimes. The issue of global definition of the frame is less important for spacetimes describing stars, which are our focus here. In the rest of this work we will fix the coordinates to be the usual $\{t,r\}$ coordinates, valid everywhere except for the irregular point $r=0$, where we need to ensure that our solutions behave in a regular manner. In these coordinates, the frame vectors are
@@
In order to solve the TOV equations, we need to specify a particular equation of state, which for a barotrope is $\mathrm{p}=\mathrm{p}(\mathrm{e})$, as well as a value for the central energy density. Then, the system of equation can be integrating outwards until $\mathrm{p}=0$, which defines the surface of the star.
@@ 5 Odd Sector
We now turn our attention to the odd sector. The perturbative variables in this case are $\{k_{u},k_{n},\beta\}$, where $k_{u,n}$ correspond to gravitational perturbations, and $\beta$ to the fluid’s angular velocity, and the only non–trivial equations are the three Einstein equations $\{\mathcal{O}_{u},\mathcal{O}_{n},\mathcal{O}_{S}\}$ and the conservation of the stress energy tensor $\mathcal{C}$. We will first review the inviscid regime, and later turn our attention to including dissipative effects.
@@ 5.1 Inviscid Regime
$\frac{\partial k_{l}}{\partial t}=\frac{e^{(\Phi-\Lambda)/2}}{2}\Bigl{(}2\frac{\partial k_{n}}{\partial r}+\frac{\partial\Phi}{\partial r}k_{n}\Bigr{)}\,.$ (34)
@@
$-\frac{\partial^{2}\psi}{\partial t^{2}}+\frac{\partial^{2}\psi}{\partial r_{\star}^{2}}-\frac{e^{\Phi}}{r^{2}}\Bigl{(}\lambda^{2}-\frac{6M}{r}+4\pi r^{2}(\mathrm{e}-\mathrm{p})\Bigr{)}\psi=0\,,$ (36)

### v2

@@ Abstract
Viscous fluids can dissipate and alter the propagation of gravitational waves, as well as modify the relaxation and stability properties of self-gravitating fluids. This is particularly relevant in order to understand the relaxation to equilibrium of neutron stars, and their gravitational wave emission. Here we study the linearized theory of perturbations of spherically symmetric self-gravitating fluids, treating only the radiative modes. Dissipative effects are included through the hydrodynamics theory of Bemfica, Disconzi, Noronha and Kovtun (BDNK). This theory has been shown to be causal and stable, despite involving only first order gradients. We show how the problem reduces to two coupled wave equations in the axial sector, one of them associated to a novel viscous mode, and including explicitly dissipative terms. In the polar sector we reduce the problem to five coupled wave equations and one additional constraint. We comment on their causal structure, and recover the causality bounds of the BDNK theory.
  • Center of Gravity, Niels Bohr Institute, Blegdamsvej 17, 2100 Copenhagen, Denmark
@@ 1 Introduction
Understanding the interaction of GWs and viscous matter also becomes important in the context of precision GW astrophysics. GW interferometers have and will observe mergers of neutron stars (NSs), which then might relax to equilibrium, or collapse promptly to form a black hole. Viscous effects can affect their GW signal during the inspiral stage Hegade K. R. et al. (2023); Ripley et al. (2023a, b); Chia et al. (2024), but also the GW emission during their final relaxation. Numerical simulations hint towards a significant impact, in particular, of bulk viscosity, in the late time relaxation of NSs Alford et al. (2010, 2018); Alford and Harris (2018); Most et al. (2021, 2022); Camelio et al. (2023a, b); Chabanov and Rezzolla (2023a, b). This happens because newly formed NSs, such as super and hypermassive NSs formed after the merger of two stars, are very hot, and bulk viscosity becomes the dominant dissipative effect at high temperatures. However, the only perturbative estimates on the relaxation timescales due to dissipative effects Cutler and Lindblom (1987); Cutler et al. (1990) are, indeed, estimates, based on the energy functional of Lindblom and Detweiler (1983). A first-principles calculation of the impact of different kinds of viscosity on the quasinormal modes of hot NSs, their precise damping times, and so on, is still lacking.
@@
In order to do this, we need to specify a particular theory of dissipative hydrodynamics. The naive extensions of the Navier–Stokes hydrodynamics to the relativistic regime, first undertaken by Eckart Eckart (1940) and Landau Landau and Lifshitz (2013) are problematic: they lead to instabilities because of their acausal character Hiscock and Lindblom (1983, 1985); Detweiler and Lindblom (1985). There is a way to heal these pathologies, by extending the theory and including effects which are second order in the gradients. This is sometimes referred to as the Israel–Stewart theory Israel and Stewart (1979); Rezzolla and Zanotti (2013), or as extended irreversible thermodynamics. However, the inclusion of second order effects complicates significantly the analysis. Recent work tackled the problem of linearized perturbations in this theory D´ıaz-Guerra et al. (2024). An alternative approach, based on a variational framework, was put forward by Carter Carter (1985, 1988, 2006); Andersson and Comer (2021). Perhaps the most natural way to solve the problems that the Eckart or Landau theories suffer from is solved simply by including all possible first order corrections to the stress energy tensor. This means including also timelike gradients, and not just spacelike ones. By following this line of thought, Bemfica, Disconzi, Noronha and Kovtun (BDNK) Bemfica et al. (2018, 2019a, 2019b, 2022); Kovtun (2019) showed that one can obtain a theory that is (i) causal, provided the transport coefficients satisfy some constraints, (ii) linearly stable, and (iii) the stress energy tensor only involves up to first order covariant derivatives of the thermodynamical variables. For this reason, in this work we will consider the stress energy tensor of a fluid in the BDNK theory as a starting point.
@@
$c_{s}$ $\sqrt{d\mathrm{p}/d\mathrm{e}}$ (Sound speed)  (17)
@@ 2 GSGM Formalism for spacetime perturbations
We use spherical symmetry to split the spacetime manifold as a warped product of a sphere and a two dimensional Lorentzian manifold, i.e., $\mathcal{M}=\mathcal{M}^{2}\times_{r^{2}}\mathcal{S}^{2}$. This decomposition allows us to reduce Einstein equations, and its perturbations, to equations on $\mathcal{M}^{2}$, by exploiting the spherical symmetry of the background. Gerlach and Sengupta first used this idea to obtain a particularly elegant derivation of the equations describing linearized perturbations around a Schwarzschild black hole Gerlach and Sengupta (1979a, 1980). Since then, this formalism has been later revisited and adapted to different set-ups, ranging from higher dimensions Kodama et al. (2000); Ishibashi and Kodama (2003); Kodama and Ishibashi (2003, 2004), further restrictions to vacuum spacetimes Martel and Poisson (2005), including deriving the equations from the curvature wave equation Mukkamala and Pereñiguez (2024), higher order perturbations Brizuela et al. (2006, 2007), perturbations in modified gravity theories Ripley and Yagi (2018); Tattersall et al. (2018), gauge theories Pereñiguez (2023), and non-stationary spacetimes Redondo-Yuste et al. (2024).
@@ 2.1 Background geometry
where $r$ is the warping factor (areal radius) and $\gamma_{ab}$ is the usual metric on the 22–sphere. Let us introduce $v_{A}=\nabla_{A}r$. Here, $\nabla$ denotes covariant derivatives on $\mathcal{M}^{2}$ compatible with the metric $g_{AB}$, and $D$ the covariant derivative on the sphere.
@@ 2.2 Linear perturbations
We now consider linearized perturbations of the metric and the stress–energy tensor. We expand them in tensor spherical harmonics following Martin-Garcia and Gundlach (1999, 2001). The tensor spherical harmonics $\{\mathcal{Z}^{\ell},\mathcal{X}^{\ell}\}$ are introduced in the Supplemental Material (SM). We also write already the metric perturbation in terms of gauge invariant variables (equivalently, in the Regge-Wheeler gauge), and discuss gauge invariance in the SM. In what follows, we consider only radiative modes, i.e., modes with spherical harmonic number $l\geq 2$. Treating the non-radiative monopolar and dipolar modes is left for future work –the expansion below in tensor spherical harmonics simplifies for $l=0,1$ modes (e.g. $\mathcal{X}^{\ell}_{a}=\mathcal{X}^{\ell}_{ab}=0$ for spherically symmetric $l=0$ modes), and residual gauge transformations can be used to further simplify their treatment. These non radiative modes do not excite propagating gravitational waves, but have important implications for the linear stability of stellar configurations.
The metric and stress energy tensor perturbations can be written as
@@
$+2\Bigl{(}k^{\ell}_{n}n_{A}-k^{\ell}_{l}l_{A}\Bigr{)}\mathcal{X}^{\ell}_{a}dy^{A}d\theta^{a}+r^{2}h^{\ell}_{S}\mathcal{Z}^{\ell}\gamma_{ab}d\theta^{a}d\theta^{b}\,$,
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.1 The general theory
$T_{\mu\nu}=\mathcal{E}u_{\mu}u_{\nu}+\mathcal{P}\perp_{\mu\nu}+u_{\mu}\mathcal{Q}_{\nu}+\mathcal{Q}_{\mu}u_{\nu}+\mathcal{T}_{\mu\nu}\,$, (16)
@@
The quantities $\mathcal{E},\mathcal{P},\mathcal{Q},\mathcal{T}$ need to be specified as functions of some thermodynamic variables. A general fluid is described by three variables: the fluid velocity $u^{\mu}$, its energy density $\mathrm{e}$ and its particle number density. In this work, in order to simplify the problem, we will restrict to barotropic matter. In particular, this corresponds to setting the heat conductivity in Ref. Bemfica et al. (2022) to zero, and neglecting the effects of particle and temperature diffusion. The heat current, however, will not vanish, as necessary to obtain a causal and stable dissipative theory. We leave for future explorations the inclusion of heat diffusion, and a chemical potential.
Now, specifying how do the terms of the stress energy tensor depend on the thermodynamic variables $\{u^{\mu},\mathrm{e}\}$ characterizes the fluid theory. For a perfect fluid, for example, one just requires that $\mathcal{E}=\mathrm{e}$, $\mathcal{P}=\mathrm{p}$ is the pressure, which is related to the energy by some equation of state, and $\mathcal{Q}=\mathcal{T}=0$. For BDNK fluids we consider additionally terms involving first order spacetime derivatives of the thermodynamic variables. In detail, the constitutive relations are Bemfica et al. (2022)
@@
where we have introduced the shorthand $\rho=\mathrm{e}+\mathrm{p}$, $\sigma_{\mu\nu}$ is the shear associated to $u_{\mu}$, and $\{\tau_{\mathrm{e}},\tau_{\mathrm{p}},\tau_{\mathcal{Q}},\eta,\zeta\}$ are the transport coefficients. Notice that the transport coefficients $\tau_{\mathrm{e},\mathrm{p},\mathcal{Q}}$ are effectively relaxation timescales, which are necessary to be added to make the theory causal. We have also introduced the sound speed $c_{s}^{2}=d\mathrm{p}/d\mathrm{e}$. We will also introduce the following combinations of transport coefficients
@@
The transport coefficients are obtained from the microscopic theory. For example, bulk viscosity in NSs is directly related to neutrino transport Camelio et al. (2023a, b). In order for the resulting hydrodynamic theory to be causal and stable, they must satisfy certain inequalities Bemfica et al. (2022). For example, in a companion work Boyanov et al. (2024) we choose a particular dependence of the transport coefficients on the stellar structure, which in turn constraints the maximum shear viscosity. However, in order to be as general as possible, here we just assume that they are smooth functions of the radial coordinate alone (in particular, we assume that they do not vary with time).
From these constitutive relations, we can immediately obtain the components of the stress energy tensor of the background. Spherical symmetry requires the fluid velocity to be $u_{\mu}dx^{\mu}=u_{A}dy^{A}$. As is natural, we choose a frame where $l_{A}=u_{A}$, and the normal vector $n_{A}=\epsilon_{AB}u^{B}$ is specified in terms of the antisymmetric symbol $\epsilon_{AB}$ associated to the background metric $g_{AB}$. Then, the constitutive relations become
@@ 3.2 Fluid perturbations
Now, in terms of these perturbed variables, we construct the stress energy tensor, and evaluate each of its components. To emphasize that we have now fixed the generic frame $\{l_{A},n_{A}\}$ in terms of the background fluid velocity, to be $\{u_{A},\epsilon_{AB}u^{B}\}$, we use the sub-index $u$ instead of $l$ to denote the contraction with the first frame vector. A straightforward but lengthy calculation yields
@@ 4 Coordinates
The previous expressions are valid for any coordinate chart (for the even parity case, for any stationary coordinate chart), which has the same domain of definition as the frame that we have introduced. We highlight that a frame is only locally defined and might not be defined, e.g., both in the exterior and in the interior region of a horizon, when dealing with black hole spacetimes. The issue of global definition of the frame is less important for spacetimes describing stars, which are our focus here. In the rest of this work we will fix the radial gauge, and denote the coordinates in this gauge by $\{t,r\}$. These are valid everywhere except for the irregular point $r=0$, where we need to ensure that our solutions behave in a regular manner. In these coordinates, the frame vectors are
@@
In order to solve the TOV equations, we need to specify a particular equation of state, which for a barotrope is $\mathrm{p}=\mathrm{p}(\mathrm{e})$, as well as a value for the central energy density. Then, the system of equation can be integrated outwards until $\mathrm{p}=0$, which defines the surface of the star.
@@ 5 Odd Sector
We now turn our attention to the odd sector. The perturbative variables in this case are $\{k_{u},k_{n},\beta\}$, where $k_{u,n}$ correspond to gravitational perturbations, and $\beta$ to the fluid’s angular velocity, and the only non–trivial equations are the three Einstein equations $\{\mathcal{O}_{u},\mathcal{O}_{n},\mathcal{O}_{S}\}$ and the conservation of the stress energy tensor $\mathcal{C}$. We will first review the inviscid regime, and later turn our attention to including dissipative effects. The odd parity sector is studied, for a particular stellar model, in the Cowling and inverse Cowling approximations in Ref. Boyanov et al. (2024). It is straightforward to obtain these approximations from the full viscous treatment, carried out here.
@@ 5.1 Inviscid Regime
$\frac{\partial k_{l}}{\partial t}=\frac{e^{(\Phi-\Lambda)/2}}{2}\Bigl{(}2\frac{\partial k_{n}}{\partial r}+\frac{d\Phi}{dr}k_{n}\Bigr{)}\,.$ (34)
@@
$\frac{\partial^{2}\psi}{\partial t^{2}}+\frac{\partial^{2}\psi}{\partial r_{\star}^{2}}-\frac{e^{\Phi}}{r^{2}}\Bigl{(}\lambda^{2}-\frac{6M}{r}+4\pi r^{2}(\mathrm{e}-\mathrm{p})\Bigr{)}\psi=0\,$, (36)